OpenStudy (anonymous):

@Mathematics f(2x + 3) = 4x + 9 f(x - 3) = 2x - 3 f(x - 3/2) = x f(-1) = 1 If f(x) = 2x + 3 and g(x) is the inverse function of f, then according to the information shown, which function represents g(x) ? A)g(x) = - 1 B)g(x) = x - 3 C)g(x) = 2x + 3 D)g(x) = x - 3/2

OpenStudy (anonymous):

f(x) = 2x+3 x = 2y + 3 y = (x-3)/2 Answer should be D

OpenStudy (anonymous):

D